# Logical Distillation of Graph Neural Networks

### v1

@@ 1 Introduction
We present and evaluate an algorithm for distilling Graph Neural Networks (GNNs) into a symbolic model. Our distillation algorithm relies on a novel model called Iterated Decision Tree (IDT), which is tailored to represent logical formulas represented by GNNs. GNNs play a crucial role in safety-critical applications like drug discovery and in cost-critical applications like large-scale transport routing. However, most GNN models are black-box in nature and their internal representations are opaque to any human or computer-aided formal scrutiny. Hence, interpreting and explaining GNN predictions is a fundamental problem of significant research interest. Although many results have characterized the expressivity of GNNs in terms of formal languages like first-order logic, extracting the logical classifiers expressed by GNNs remains largely unexplored. We aim to fill this gap by developing a distillation model aimed at extracting logical classifiers expressed by GNNs.
The key theoretical insight behind our model is that all the first-order logic classifiers expressed by GNNs are expressible in first-order logic with only two variables and counting quantifiers ($\mathrm{C}^{2}$). Hence our model, the IDT, is designed to express any $\mathrm{C}^{2}$ formula. An IDT consists of a sequence of decision trees. Each decision tree expresses a number of unary $\mathrm{C}^{2}$ formulas of quantifier depth one. Combining multiple such decision trees enables us to express formulas of larger quantifier depth. Additionally, we propose an extension of $\mathrm{C}^{2}$ that can capture operations like mean aggregation, which are common in GNNs, and incorporate it into IDTs. Our distillation algorithm is able to exploit intermediate node representations from each message-passing layer of a GNN to iteratively learn decision trees of an IDT. Although the learning process for IDTs is guided by the GNN, our empirical results show that the logic-based inductive bias incentivizes succinct and interpretable models.
@@ 2 Related Work
Our work is also loosely connected to the general problem of learning desision trees from neural networks. This problem has already been extensively investigated for tabular data (Craven and Shavlik, 1995; Krishnan, Sivakumar, and Bhattacharya, 1999; Boz, 2002; Dancey, McLean, and Bandar, 2004; Setzu et al., 2021). Furthermore, recent works have also investigated the tweaking of learning process or the neural architecture itself for learning decision trees (Schaaf, Huber, and Maucher, 2019; Wu et al., 2018; Yang, Morillo, and Hospedales, 2018; Kontschieder et al., 2016). Although these results are related to our approach in spirit, our work is fundamentally different in its theoretical motivation and the learning procedure. GNNs expressivity is deeply connected to that of first-order logic (Barceló et al., 2020; Grohe, 2023). Hence, using logic-based decision trees is a natural choice for learning decision trees from GNNs.
Explanation methods that distill surrogate models GNNs come closest to our approach. Azzolin et al. (2023) first derive instance-level local subgraphs as explanations and then cluster them to extrapolate a model-level Boolean formula using the subgraphs as concepts. Yuan et al. (2020) base their approach on input-optimization, i.e. globally reducing graphs to a number of instances for which the explanation is then given. Their approach requires prior domain knowledge. Most recently, Müller et al. (2024) first compute (almost) categorical layer wise node representations using GNN layers with Gumbel-Softmax update functions. Subsequently, they replace the neural networks by decision trees trained on the categorical node representations. Their approach results in an interpretable message passing scheme based on intermediate categorical node states, but requires to train a specific GNN architecture. All three approaches use graphs, sub-graphs or their combinations as the explanation model. This restricts these methods, as many simple and important constraints, e.g. a graph has more than 12 nodes, can not easily be expressed in terms of subgraphs.
@@ 3.2 Graphs and Logic
Theorem 1 (Barceló et al. (2020)).
A classifier that is expressible in first-order logic is expressible in $\mathrm{C}^{2}$ if and only if it can be computed by a graph neural network.
Note that Theorem 1 completely characterizes the first-order logic expressivity of GNNs. However, GNNs can also express classifiers that are inexpressible in first-order logic. The proof of Theorem 1 builds on the equivalence between $\mathrm{C}^{2}$ and the modal logic $\mathcal{EMLC}$ (Lutz, Sattler, and Wolter, 2001). $\mathcal{EMLC}$ allows us to define a convenient grammar and serves as the motivation for our proposed model. We follow Barceló et al. (2020, Appendix D) and introduce a language similar to their Lemma D.4.
@@ 6 Relative Modal Parameters
holds for if more than half the nodes in $G$ satisfy $U_{0}$.
@@ 8 Conclusion
We have presented Iterated Decision Trees (IDTs), a novel model tailored towards distilling interpretable logical formulas from Graph Neural Networks (GNNs). Building on previous results, we have shown that IDTs can express every first-order logic formula that can be captured by a GNN. We have also introduced an extension of the fragment of first-order logic that can be captured by GNNs. This extension was easily incorporated into IDTs without any significant computational overhead. The distilled IDTs often surpass the accuracy of the underlying GNN while providing insight into the decision process. They also outperform the considered GNNs when the ground truth is itself a logical formula. The classification decisions of the IDT are interpretable and enable us to extract insights on multiple datasets.

### v2

@@ 1 Introduction
We present and evaluate an algorithm for distilling Graph Neural Networks (GNNs) into a symbolic model. Our distillation algorithm relies on a novel model called Iterated Decision Tree (IDT), which is tailored to represent logical formulas represented by GNNs. GNNs play a crucial role in safety-critical applications like drug discovery and in cost-critical applications like large-scale transport routing. However, most GNN models are black-box in nature and their internal representations are opaque to any human or computer-aided formal scrutiny. Hence, interpreting and explaining GNN predictions is a fundamental problem of significant research interest. Although many results have analyzed the expressivity of GNNs in terms of formal languages like first-order logic, extracting the logical classifiers expressed by GNNs remains largely unexplored. We aim to fill this gap by developing a distillation model aimed at extracting logical classifiers expressed by GNNs.
The key motivation for our model is the close relationship between GNNs and first-order logic with only two variables and counting quantifiers $\mathrm{C}^{2}$ (Barceló et al., 2020; Grohe, 2021, 2023). Hence our model, the IDT, is designed to express any $\mathrm{C}^{2}$ formula. An IDT consists of a sequence of decision trees. Each decision tree expresses a number of unary $\mathrm{C}^{2}$ formulas of quantifier depth one. Combining multiple such decision trees enables us to express formulas of larger quantifier depth. Additionally, we propose an extension of $\mathrm{C}^{2}$ that can capture operations like mean aggregation, which are common in GNNs, and incorporate it into IDTs. Our distillation algorithm is able to exploit intermediate node representations from each message-passing layer of a GNN to iteratively learn decision trees of an IDT. Although the learning process for IDTs is guided by the GNN, our empirical results show that the logic-based inductive bias incentivizes succinct and interpretable models.
@@ 2 Related Work
Our work is also loosely connected to the general problem of learning desision trees from neural networks. This problem has already been extensively investigated for tabular data (Craven and Shavlik, 1995; Krishnan, Sivakumar, and Bhattacharya, 1999; Boz, 2002; Dancey, McLean, and Bandar, 2004; Setzu et al., 2021). Furthermore, recent works have also investigated the tweaking of learning process or the neural architecture itself for learning decision trees (Schaaf, Huber, and Maucher, 2019; Wu et al., 2018; Yang, Morillo, and Hospedales, 2018; Kontschieder et al., 2016). Although these results are related to our approach in spirit, our work is fundamentally different in its theoretical motivation and the learning procedure. GNNs expressivity is deeply connected to that of first-order logic (Barceló et al., 2020; Grohe, 2021, 2023). Hence, using logic-based decision trees is a natural choice for learning decision trees from GNNs.
Explanation methods that distill surrogate models from GNNs come closest to our approach. Azzolin et al. (2023) first derive instance-level local subgraphs as explanations and then cluster them to extrapolate a model-level Boolean formula using the subgraphs as concepts. Yuan et al. (2020) base their approach on input-optimization, i.e. globally reducing graphs to a number of instances for which the explanation is then given. Their approach requires prior domain knowledge. Most recently, Müller et al. (2024) first compute (almost) categorical layer wise node representations using GNN layers with Gumbel-Softmax update functions. Subsequently, they replace the neural networks by decision trees trained on the categorical node representations. Their approach results in an interpretable message passing scheme based on intermediate categorical node states, but requires to train a specific GNN architecture. All three approaches use graphs, sub-graphs or their combinations as the explanation model. This restricts these methods, as many simple and important constraints, e.g. a graph has more than 12 nodes, can not easily be expressed in terms of subgraphs.
@@ 3.2 Graphs and Logic
Theorem 1 (Barceló et al. (2020), Theorem 5.2).
Any classifier expressible in $\mathrm{C}^{2}$ can be computed by a GNN.
Note that GNNs can also compute classifiers that are inexpressible in first-order logic. Whether every first-order classifier computed by a GNN is expressible in $\mathrm{C}^{2}$ is an open question. However, for GNNs without $\mathsf{pool}$ operation, first-order logic expressivity is completely characterized by the guarded fragment of $\mathrm{C}^{2}$ (Barceló et al., 2020, Theorem 4.2). The proof of Theorem 1 builds on the equivalence between $\mathrm{C}^{2}$ and the modal logic $\mathcal{EMLC}$ (Lutz, Sattler, and Wolter, 2001). $\mathcal{EMLC}$ allows us to define a convenient grammar and serves as the motivation for our proposed model. We follow Barceló et al. (2020, Appendix D) and introduce a language similar to their Lemma D.4.
@@ 6 Relative Modal Parameters
holds if more than half the nodes in $G$ satisfy $U_{0}$.
@@ 8 Conclusion
We have presented Iterated Decision Trees (IDTs), a novel model tailored towards distilling interpretable logical formulas from Graph Neural Networks (GNNs). IDTs can express any logical formula expressible in first order logic with two variables and counting quantifiers ($\mathrm{C}^{2}$) — a fragment of first-order logic that is closely connected to the logical expressivity of GNNs. We have also introduced an extension of $\mathrm{C}^{2}$ that captures operations commonly used in GNNs. This extension was easily incorporated into IDTs without any significant computational overhead. The distilled IDTs often surpass the accuracy of the underlying GNN while providing insight into the decision process. They also outperform the considered GNNs when the ground truth is itself a logical formula. The classification decisions of the IDT are interpretable and enable us to extract insights on multiple datasets.